# Survey on Large Language Model-Enhanced Reinforcement Learning: Concept, Taxonomy, and Methods

### v1

@@ Abstract
With extensive pre-trained knowledge and high-level general capabilities, large language models (LLMs) emerge as a promising avenue to augment reinforcement learning (RL) in aspects such as multi-task learning, sample efficiency, and task planning. In this survey, we provide a comprehensive review of the existing literature in LLM-enhanced RL and summarize its characteristics compared to conventional RL methods, aiming to clarify the research scope and directions for future studies. Utilizing the classical agent-environment interaction paradigm, we propose a structured taxonomy to systematically categorize LLMs’ functionalities in RL, including four roles: information processor, reward designer, decision-maker, and generator. Additionally, for each role, we summarize the methodologies, analyze the specific RL challenges that are mitigated, and provide insights into future directions. Lastly, potential applications, prospective opportunities and challenges of the LLM-enhanced RL are discussed.
@@
Reinforcement learning (RL), large language models (LLM), vision-language models (VLM), multi-modal RL, LLM-enhanced RL.
@@ I Introduction
Reinforcement learning (RL) is a powerful learning paradigm that focuses on control and decision-making, where an agent learns to optimize a specified target through trial and error interactions with the environment. The recent success of deep learning in classical domains, such as natural language processing (NLP) [1] and computer vision (CV) [2], has led to the rise of deep RL, which combines RL with high-capacity deep neural network approximators that enable agents to make decisions in complex environments. In the game realm, deep RL algorithms have achieved the beyond-human ability of decision-making in a multitude of arcade [3][4], real-time strategy [5][6], board [7][8], and imperfect information games [9][10]. Currently, the application of RL has to extend to real-world tasks in the fields of robotics [11], autonomous vehicles [12], healthcare [13], and conversational system [14].
@@
However, deep RL algorithms face significant challenges when applied in real-world applications involving language and visual information, as the agent has to learn the features and a control policy jointly. To reduce the burden of visual feature learning, reference [15] decouples representation learning from RL. To handle language-involved tasks, a survey [16] calls for potential uses of NLP techniques in RL. Nevertheless, the capabilities of language models were limited at that time and the following four challenges still have not been addressed: 1) Sample inefficiency: deep RL agents require extensive interactions with the environment to learn effective policies, rendering them impractical for data collection scenarios that are expensive or risky [17][18][19]. 2) Reward function design: the performance of policy learning significantly depends on the design of the reward function [20]. Despite the fundamental importance, reward functions are known to be notoriously difficult to design [21][22] since they require a deep understanding of the task and often demand manual trial-and-error [23]. 3) Generalization: the generalization of deep RL agents remains daunting, as they frequently struggle to adapt to new, unseen environments, limiting the agents’ applicability in dynamic real-world settings [24][25]. 4) Natural language understanding: deep RL faces difficulties in natural language processing and understanding scenarios, where the nuances and complexities of human language present unique challenges that are not adequately addressed by current RL methodologies [26].
@@
The recent emergence of large language models (LLMs) has marked a significant milestone in the field of NLP and shown various powerful capabilities in many real-world applications such as medicine [27], chemical [28] and embodied control in robots [29]. Compared to small language models, LLMs have emergent capabilities that are not present in small language models [30], such as in-context learning [31], reasoning ability [32] etc. Additionally, leveraging the vast amounts of training data, pre-trained LLMs are equipped with a broad spectrum of world knowledge [33]. Benefiting from these capabilities, the applications of language models have been shifted from language modeling to task-solving, ranging from basic text classification and sentiment analysis to complex high-level task planning [34] and decision-making [35][36][37].
@@
With emergent capabilities, the potential of LLMs to address the inherent challenges of RL has recently gained popularity [38][39]. The capabilities of LLMs, particularly in natural language understanding, reasoning, and task planning, provide a unique approach to solving the above-mentioned RL issues. For sample inefficiency, reference [40] proposes a framework where LLMs can be employed to improve the sample efficiency of RL agents by providing rich, contextually informed predictions or suggestions, thereby reducing the need for extensive environment interactions. For reward function design, LLMs can aid in constructing more nuanced and effective reward functions, enhancing the learning process by offering a deeper understanding of complex scenarios [41]. For generalization, reference [42] proposes a framework that leverages language-based feedback for improving the generalization of RL policy in unseen environments. For natural language understanding, Pang et al.  [43] use LLMs to translate complex natural language-based instructions to simple task-specified languages for RL agents. These works show that LLM is a promising and powerful role that can contribute to the longstanding RL challenges.
@@ II Background
In this section, we provide a concise overview of the classical RL paradigm and related challenges. Next, we explore the prevailing trend in RL—specifically, the fusion of multi-modal data sources, including language and visual information. Following this, we offer an introductory background on LLMs and outline the key capabilities that can enhance the RL learning paradigm.
@@ II-A2 Challenges of Reinforcement Learning
II-A3 Multi-modal Reinforcement Learning
With the advances in both CV and NLP, pattern recognition in vision and natural language has become increasingly powerful, and multi-modal data has been involved in the RL paradigm recently. Visual data is commonly involved in the observation space of RL when agents receive image-based information from the environment, e.g. in applications such as robots [51], video game control [3] etc. Compared to visual data, natural languages are usually included when RL agents are given specific tasks when interacting with the environments. The use of natural languages in RL can be divided into the following two categories [16]:
@@ II-A3 Multi-modal Reinforcement Learning
The integration of multi-modal data challenges the RL paradigm since the agent has to simultaneously learn how to process complex multi-modal data and optimize the control policy in the environment [15]. Issues such as natural language understanding [52] and visual-based reward function design [53] require to be addressed.
@@ II-B Background of Large Language Models
LLMs typically refer to the Transformer-based language models [54] containing hundreds of billions of parameters and being trained on massive text data (i.e., several terabytes (TB) scale) [55], such as GPT-3 [56], Galactica [57], PaLM [58] and LLaMA [59]. The vast number of parameters and internet-scale training data enable LLMs to learn a wide range of tasks, enabling more powerful abilities in language generation, knowledge representation, logical reasoning, etc., and better generalization capabilities on new tasks.
@@
  • Instruction Following: Leveraging diverse task-specific datasets formatted with natural language descriptions (also called instruction tuning), LLMs are shown to perform well on unseen tasks that are also described in the form of natural language [63][64]. Therefore, this capability equips LLMs with the ability to comprehend instructions for new tasks and effectively generalize across tasks not previously encountered, even in the absence of explicit examples.
@@ III-A Definition
RL agents are often tasked with making robust and deliberate decisions using multi-modal information in real-real applications, whether in the MDP setting or within the context of specific task descriptions. Examples include robots designed to follow natural language instructions while navigating physical environments or visual games with tasks described in natural language [68][69][70]. However, it is challenging for conventional RL methods as the agent is required to simultaneously interpret complex multi-modal data and optimize control policies amidst ever-changing environments [71]. Compounding these challenges are issues like sample inefficiency, the difficulty of crafting reward functions that accurately reflect multi-modal inputs, and the need for robust generalization across varied tasks and settings.
The rapid advancements in LLMs present a viable solution to these challenges, thanks to their potent natural language understanding and reasoning abilities, coupled with recent progress in incorporating visual data processing [72]. This dual capability enables LLMs to interpret and act upon complex multi-modal information effectively, serving as a robust helper for enhancing the RL paradigm for real-world applications.
@@
LLM-enhanced RL refers to the methods that utilize the multi-modal information processing, generating, reasoning, etc. capabilities of pre-trained, knowledge-inherent AI models to assist the RL paradigm.
Unlike traditional model-based reinforcement learning, a key characteristic of LLM-enhanced RL is that it utilizes the models with general knowledge, which means that the model has a considerable level of ability at the beginning of the learning process and better generalization compared to other data-driven models. In addition, the LLM-enhanced RL still uses the AI model and can continuously learn through the RL interaction data to improve its capabilities.
@@ III-C Characteristics
  • Multi-modal Information Understanding: LLMs enhance RL agents’ comprehension of scenarios involving multi-modal information, enabling them to learn from tasks or environments described in natural language and vision data more effectively.
  • Multi-task Learning and Generalization: Benefiting from the multi-disciplinary pre-trained knowledge, LLMs empower RL agents by either designing rewards or generating reference actions, facilitating multi-task learning and generalization.
@@ III-D Taxonomy
  • Information Processor: When observation or task description involves language or visual features, it is challenging for the agent to comprehend the complex information and optimize the control policy simultaneously. To release the agent from the burden of understanding the multi-modal data, LLM can serve as an information processor for environment information or task instruction information by 1): extracting meaningful feature representations for speeding up network learning; 2) translating natural language-based environment information or task instruction information into formal specific task languages to reduce learning complexity.
@@
  • Decision-maker: RL faces the challenges of sample inefficiency and planning when exploring long-horizon environments. By serving as the decision-maker, pre-trained LLMs can perform: 1) direct decision-making: use the powerful sequence modeling ability and common-sense knowledge to improve the sample efficiency of offline RL; 2) indirect decision-making: act as an expert instructor that generates action candidates (high-level task planning) to shrink the action selection range or outputs reference policies to indirectly guide the updating directions of RL policies.
  • Generator: Model-based RL hinges on precise world models to learn accurate environment dynamics and simulate high-fidelity trajectories. Additionally, interpretability remains another important issue in RL. Using the multi-modal information understanding capability and prior common-sense reasoning ability, LLMs can be 1) a generator to generate accurate trajectories in model-based RL; 2) generate policy explanations with the prompts of related information in explainable RL.
@@ IV LLM as Information Processor
With the advances in unsupervised techniques and large-scale pre-trained models for CV and NLP, the decoupled structure, where the encoders are separately trained, has gained popularity [15][73][74]. Utilizing the powerful representation ability and prior knowledge, the pre-trained LLM or vision-language model (VLM) model can serve as an information processor to either extract an observation representation for downstream networks or translate the unbounded natural language to formal specification languages, which increase the sample efficiency and zero-shot performance of RL.
@@ IV-A Feature Representation Extractor
Figure 3: LLM as an information processor.
@@ IV-A1 Frozen Pre-trained Model
Using the frozen large-scale pre-trained model is the straightforward way. In reference [75], the author proposes History Compression via Language Models (HELM) to utilize a frozen pre-trained Language Transformer to extract history representation and compression and thus addresses the problem of partially observed MDP by approximating the underlying MDP with the past representation. Specifically, the framework first uses FrozenHopfield, a frozen associative memory to map observations $[o_{t-2},o_{t-1},o_{t}]$ to a compressed representation $h_{t}$ and then concatenate it with a learned encoding of the current observation via a convolutional neural network. Such a method solves the problem of how to effectively utilize the compressed history for policy optimization. After that, Semantic HELM [76] proposes a human-readable memory mechanism that summarizes past visual observations in human language and uses multimodal models to associate visual inputs with language tokens. The memory is a semantic database $\mathcal{S}$ constructed by encoding prompt-augmented tokens from the vocabularies of Contrastive Language-Image Pre-training (CLIP) [77] and the pre-trained language models. Given an observation $o_{t}$, the agent retrieves top-$k$ embeddings from $\mathcal{S}$ as actor-critic input to assist the policy optimization. Such a memory mechanism provides a human-readable representation of the past and helps the agent to cope with partially observable environments. However, one limitation of such frozen pre-trained models is that the representations cannot dynamically adjust according to the task and environment.

### v3

@@ Abstract
With extensive pre-trained knowledge and high-level general capabilities, large language models (LLMs) emerge as a promising avenue to augment reinforcement learning (RL) in aspects such as multi-task learning, sample efficiency, and high-level task planning. In this survey, we provide a comprehensive review of the existing literature in LLM-enhanced RL and summarize its characteristics compared to conventional RL methods, aiming to clarify the research scope and directions for future studies. Utilizing the classical agent-environment interaction paradigm, we propose a structured taxonomy to systematically categorize LLMs’ functionalities in RL, including four roles: information processor, reward designer, decision-maker, and generator. For each role, we summarize the methodologies, analyze the specific RL challenges that are mitigated, and provide insights into future directions. Lastly, comparative analysis of each role, potential applications, prospective opportunities and challenges of the LLM-enhanced RL are discussed. By proposing this taxonomy, we aim to provide a framework for researchers to effectively leverage LLMs in the RL field, potentially accelerating RL applications in complex applications such as robotics, autonomous driving, and energy systems.
@@
Reinforcement learning (RL), large language models (LLM), vision-language models (VLM), multimodal RL, LLM-enhanced RL.
@@ I Introduction
Reinforcement learning (RL) is a powerful learning paradigm that focuses on control and decision-making, where an agent learns to optimize a specified target through trial and error interactions with the environment. Traditional RL methods, however, often struggled with high-dimensional state spaces and complex environments [1]. The integration of deep learning techniques with RL, known as deep RL, has led to significant breakthroughs. In 2015, Deep Q-Networks (DQN) [2] marked a turning point, demonstrating human-level performance on Atari games using raw pixel inputs. Subsequent innovations such as proximal policy optimization [3] and soft actor-critic [4] have further expanded the capabilities of deep RL. In different realms, deep RL algorithms have achieved promising performance, such as real-time strategy games [5][6], board games [7][8], energy management [9] and imperfect information games [10][11]. Concurrent advancements in natural language processing (NLP) and computer vision (CV) [12, 13], have fostered new RL paradigms, such as language-conditional RL [14], which uses natural language to instruct agents, and vision-based RL [15], where agents learn from high-dimensional visual inputs.
@@
The integration of language and vision capabilities into deep RL has introduced new challenges, as the agent has to learn the high-dimensional features and a control policy jointly. To reduce the burden of visual feature learning, reference [16] decoupled representation learning from RL. To handle language-involved tasks, a survey [17] called for potential uses of NLP techniques in RL. Nevertheless, the capabilities of language models were limited at that time and the following four challenges still have not been addressed: 1) Sample inefficiency: Language and vision tasks involve large, complex state-action spaces, making it challenging for RL agents to learn effective policies. Moreover, agents must understand tasks and connect them to corresponding states, necessitating even more extensive interactions [18][19][20]. 2) Reward function design: In language and vision tasks, designing effective reward functions is particularly challenging. These functions must capture subtle linguistic nuances and complex visual features, significantly increasing the complexity of an already difficult process. Moreover, aligning rewards with high-level task objectives in these domains often requires domain expertise and extensive trial-and-error [21][22][1][23]. 3) Generalization: RL agents often overfit to training data, especially in vision-based environments, leading to poor performance when deployed in states with interventions (e.g., added noise). Agents must learn invariant features robust to such interventions, enabling generalization across varied linguistic contexts and visual scenes. However, the complexity of these domains makes extracting such features and adapting to new environments particularly challenging [24][25]. 4) Natural language understanding: deep RL faces difficulties in natural language processing and understanding scenarios, where the nuances and complexities of human language present unique challenges that are not adequately addressed by current RL methodologies [26].
@@
The field of NLP has undergone a revolutionary transformation since the introduction of the Transformer architecture in 2017 [12]. This breakthrough paved the way for the development of Large Language Models (LLMs), with landmark models such as BERT [27], GPT [28], and more recent iterations such as GPT-3 [29] and PaLM [30] marking significant milestones. The emergence of these LLMs has demonstrated powerful capabilities across various real-world applications, including medicine [31], chemical research [32], energy system [33][34][35] and embodied control in robotics [36]. These models have not only advanced the field of NLP but also shown remarkable potential in tackling complex, multi-disciplinary challenges. Compared to small language models, LLMs have emergent capabilities that are not present in small language models [37], such as in-context learning [38], reasoning ability [39] etc. Additionally, leveraging the vast amounts of training data, pre-trained LLMs are equipped with a broad spectrum of world knowledge [40]. Benefiting from these capabilities, the applications of language models have been shifted from language modeling to task-solving, ranging from basic text classification and sentiment analysis to complex high-level task planning [41] and decision-making [42][43].
@@
With emergent capabilities, the potential of LLMs to address the inherent challenges of RL has recently gained popularity [44][45]. The capabilities of LLMs, particularly in natural language understanding, reasoning, and task planning, provide a unique approach to solving the above-mentioned RL issues. For sample inefficiency, reference [46] proposed a framework where LLMs can be employed to improve the sample efficiency of RL agents by providing rich, contextually informed predictions or suggestions, thereby reducing the need for extensive environment interactions. For reward function design, LLMs can aid in constructing more nuanced and effective reward functions, enhancing the learning process by offering a deeper understanding of complex scenarios [47]. For generalization, reference [48] proposed a framework that leverages language-based feedback for improving the generalization of RL policy in unseen environments. For natural language understanding, Pang et al.  [49] used LLMs to translate complex natural language-based instructions to simple task-specified languages for RL agents. These works show that LLM is a promising and powerful role that can contribute to the longstanding RL challenges.
@@ II Background
In this section, we provide a concise overview of the classical RL paradigm and related challenges. Next, we explore the prevailing trend in RL—specifically, the fusion of multimodal data sources, including language and visual information. Following this, we offer an introductory background on LLMs and outline the key capabilities that can enhance the RL learning paradigm.
@@ II-A2 Challenges of Reinforcement Learning
  • Multi-task Learning: Multi-task RL faces several key challenges that limit its effectiveness. One major issue is managing varying task difficulties, where simpler tasks can overshadow learning of more complex ones, leading to negative transfer [57]. Task interference is another critical problem, as shared parameters or data between tasks can result in suboptimal performance on individual tasks [58]. Determining optimal parameter-sharing strategies is complex, as it must balance learning efficiency with task-specific requirements [59]. Sample efficiency remains a significant hurdle, with traditional data-sharing approaches not fully leveraging learned behaviors across tasks [60]. Finally, effectively transferring knowledge between tasks without negative interference is an ongoing challenge that impacts the agent’s ability to accelerate learning across multiple objectives [61].
II-A3 Multimodal Reinforcement Learning
With the advances in both CV and NLP, pattern recognition in vision and natural language has become increasingly powerful, and multimodal data has been involved in the RL paradigm recently. Visual data is commonly involved in the observation space of RL when agents receive image-based information from the environment, e.g. in applications such as robots [62], video game control [2] etc. Compared to visual data, natural languages are usually included when RL agents are given specific tasks when interacting with the environments. The use of natural languages in RL can be divided into the following two categories [17]:
@@ II-A3 Multimodal Reinforcement Learning
The integration of multimodal data challenges the RL paradigm since the agent has to simultaneously learn how to process complex multimodal data and optimize the control policy in the environment [16]. Issues such as natural language understanding [63] and visual-based reward function design [64] require to be addressed.
@@ II-B Background of Large Language Models
LLMs typically refer to the Transformer-based language models [65] containing billions of parameters and being trained on massive text data (i.e., several terabytes (TB) scale) [66], such as GPT-3 [29] and LLaMA [67]. The extensive number of parameters and internet-scale training data enable LLMs to master a diverse array of tasks, resulting in enhanced capabilities in language generation, knowledge representation, and logical reasoning, as well as improved generalization to novel tasks.
@@
  • Instruction Following: Leveraging diverse task-specific datasets formatted with natural language descriptions (also called instruction tuning), LLMs are shown to perform well on unseen tasks that are also described in the form of natural language [71][72][73]. Therefore, this capability equips LLMs with the ability to comprehend instructions for new tasks and effectively generalize across tasks not previously encountered, even in the absence of explicit examples.
@@ III-A Definition
RL agents are often tasked with making robust and deliberate decisions using multimodal information in real-real applications, whether in the MDP setting or within the context of specific task descriptions. Examples include robots designed to follow natural language instructions while navigating physical environments or visual games with tasks described in natural language [77][78][79]. However, it is challenging for conventional RL methods as the agent is required to simultaneously interpret complex multimodal data and optimize control policies amidst ever-changing environments [80]. Compounding these challenges are issues like sample inefficiency, the difficulty of crafting reward functions that accurately reflect multimodal inputs, and the need for robust generalization across varied tasks and settings.
The rapid advancements in LLMs present a viable solution to these challenges, thanks to their potent natural language understanding and reasoning abilities, coupled with recent progress in incorporating visual data processing [81]. This dual capability enables LLMs to interpret and act upon complex multimodal information effectively, serving as a robust helper for enhancing the RL paradigm for real-world applications.
@@
LLM-enhanced RL refers to the methods that utilize the multimodal information processing, generating, reasoning, and other high-level cognitive capabilities of pre-trained, knowledge-inherent LLM models to assist the RL paradigm.
LLM-enhanced RL differs from traditional model-based RL by leveraging knowledge-rich LLM models. This approach provides two key advantages: First, LLM equips the agent with substantial pre-trained capabilities at the beginning of the learning process, such as reasoning and high-level planning, etc. Second, it offers superior generalization capabilities. Pre-trained on diverse data, LLMs can effectively transfer knowledge across domains, enabling better adaptation to unseen environments than conventional data-driven models. Last, LLM-enhanced RL addresses a key limitation of pre-trained models: their inability to interact with environments to expand their knowledge and ground themselves in specific domains. Through environmental interactions, this approach generates task-specific data, grounds the LLM in particular domains by in-context learning, and eventually helps them adapt to dynamic changes with continuous learning.
@@ III-C Characteristics
  • Multimodal Information Understanding: LLMs enhance RL agents’ comprehension of scenarios involving multimodal information, enabling them to learn from tasks or environments described in natural language and vision data more effectively.
  • Multi-task Learning and Generalization: Benefiting from the multi-disciplinary pre-trained knowledge and powerful sequence modeling capability, LLMs empower RL agents by providing a high-capacity model capable of accommodating task variances and transferring knowledge across multiple tasks, assisting in handling multiple tasks.
@@ III-D Taxonomy
  • Information Processor: When observation or task description involves language or visual features, it is challenging for the agent to comprehend the complex information and optimize the control policy simultaneously. To release the agent from the burden of understanding the multimodal data, LLM can serve as an information processor for environment information or task instruction information by 1): extracting meaningful feature representations for speeding up network learning; 2) translating natural language-based environment information or task instruction information into formal specific task languages to reduce learning complexity.
    Application example: In instruction-following RL for robots, tasks can have unbounded natural language forms due to users’ diverse speaking habits, which can impede RL learning performance. LLMs transform these varied natural language instructions into a unique task language, enabling more robust RL performance [49].
@@
    Application example: For complex robotic control problems, such as dexterous manipulation, reward design requires both expertise and trial and error. LLM designs reward functions based on knowledge and iteratively improves it based on the performance [82].
  • Decision-maker: RL faces challenges such as sample inefficiency and exploration inefficiency. To address these challenges, LLMs can be leveraged as decision-makers in RL, offering promising solutions through two main approaches: 1) action-making: LLMs treat offline RL as a sequence modeling problem, using rewards for the conditional generation of actions. Pretrained on diverse, internet-scale data, LLMs possess advanced semantic understanding capabilities, which can be exploited to accelerate offline RL learning. 2) action-guiding: LLMs act as expert instructors to produce a reduced set of action candidates or expert actions. The action candidates constrain the original action space, thereby enhancing exploration efficiency. Expert actions encapsulate prior knowledge from LLMs. When incorporated to regularize the policy learning, this expert knowledge is distilled into an RL agent, resulting in better sample efficiency.
    Application example: In embodied robot, given human-instruction and language-based description, LLM generates potential actions for the robot to choose from [83].
  • Generator: Model-based RL hinges on precise world models to learn accurate environment dynamics and simulate high-fidelity trajectories. Additionally, interpretability remains another important issue in RL. Using the multimodal information understanding capability and prior common-sense reasoning ability, LLMs can be 1) a generator to generate accurate trajectories in model-based RL; 2) generate policy explanations with the prompts of related information in explainable RL. Application example: In Minecraft item crafting, LLMs generate Abstract World Models—hypothesized sequences of subgoals for a given task. These LLM-generated world models guide RL agents’ exploration and learning and the RL agent verifies and corrects the world model through gameplay, combining LLM knowledge with grounded experience to achieve an order of magnitude improvement in sample efficiency over traditional methods [84].
@@ IV LLM as Information Processor
With the advances in unsupervised techniques and large-scale pre-trained models for CV and NLP, the decoupled structure, where the encoders are separately trained, has gained popularity [16][85][86]. Utilizing the powerful representation ability and prior knowledge, the pre-trained LLM or vision-language model (VLM) model can serve as an information processor for RL. They can extract observation representations for downstream networks or translate natural language into formal specifications, enabling the execution of multiple tasks. This multi-task capability improves sample efficiency and zero-shot performance, allowing agents to generalize effectively across diverse and sparse-reward environments.
@@ IV-A Feature Representation Extractor
Figure 3: LLM as an information processor. (i) Feature Representation Extractor: frozen/fine-tuned LLM extracts meaningful representations for downstream RL networks. In the fine-tuning process, given observation ($\mathcal{O}_{t}$), invariant feature abstraction ($\tilde{S}_{t}$) is learned with the contrastive loss ($\mathcal{L}^{c}_{t}$). Then, the invariant is fed into the actor-critic network. After fine-tuning, given different observations ($O_{t}$) and ($O^{\prime}_{t}$) with appearance variation, the extracted representation is invariant, leading to robust RL performance. (ii) Language Translator: LLM interprets diverse natural language inputs, converting them into a standardized, task-specific format that the RL agent can efficiently process and act upon.
@@ IV-A1 Frozen Pre-trained Model
Using the frozen large-scale pre-trained model is the straightforward way. In reference [87], the author proposed History Compression via Language Models (HELM) to utilize a frozen pre-trained Language Transformer to extract history representation and compression and thus addresses the problem of partially observed MDP by approximating the underlying MDP with the past representation. Specifically, the framework first uses FrozenHopfield, a frozen associative memory to map observations $[o_{t-2},o_{t-1},o_{t}]$ to a compressed representation $h_{t}$ and then concatenate it with a learned encoding of the current observation via a convolutional neural network. Such a method solves the problem of how to effectively utilize the compressed history for policy optimization. After that, Semantic HELM [88] proposed a human-readable memory mechanism that summarizes past visual observations in human language and uses multimodal models to associate visual inputs with language tokens. The memory is a semantic database $\mathcal{S}$ constructed by encoding prompt-augmented tokens from the vocabularies of Contrastive Language-Image Pre-training (CLIP) [89] and the pre-trained language models. Given an observation $o_{t}$, the agent retrieves top-$k$ embeddings from $\mathcal{S}$ as actor-critic input to assist the policy optimization. Such a memory mechanism provides a human-readable representation of the past and helps the agent to cope with partially observable environments. In the experiment, they used the proximal policy optimization (PPO) [3] algorithm and pretrained Transformer XL [90] model. When testing on the partially observable environments, they found the extracted semantics help the memory-less agent obtain comparable scores with memory-based methods trained on long trajectories. However, one limitation of such frozen pre-trained models is that the representations cannot dynamically adjust according to the task and environment.